Complete these structures by adding bonds and lone pairs as necessary. Then, determine the nitrogen-to-oxygen bond order in each ion.

Complete these structures by adding bonds and lone
Answer #1
Concepts and reason

• In Chemistry, Lewis structure represents the number of electrons present around each atom of molecule as a bonding electrons as well as non-bonding electrons. For molecule, the number of bonds and lone pairs can be added following the octet rule.

• Bond order of the molecule represents the number of bonds present between the atoms. Bond order is the measurement of the bond strength of a molecule. An increase in the bond order shows the increase in the bond strength.

Fundamentals

Bond order is calculated using the following formula:

Bondorder=TotalnumberofelectronpairsinNObondTotalnumberofNObondgroups{\rm{Bond}}\,{\rm{order}}\,{\rm{ = }}\,\frac{{{\rm{Total number of electron pairs in N - O bond}}}}{{{\rm{Total number of}}\,{\rm{N - O bond groups}}}}

The structure of NO3{\rm{NO}}_3^ - molecule is given below:

NO;

From the given structure of NO3{\rm{NO}}_3^ - molecule,

The total number of electron pairs in NO{\rm{N}} - {\rm{O}} bond = 4 electron pairs.

The total number of NO{\rm{N}} - {\rm{O}} bond groups = 3bonds (1 double bond, 2 single bonds).

The bond order is calculated below:

Bondorder=TotalnumberofelectronpairsinNObondTotalnumberofNObondgroupsBondorder=43=1.3.\begin{array}{c}\\{\rm{Bond}}\,{\rm{order}}\,{\rm{ = }}\,\frac{{{\rm{Total number of electron pairs in N - O bond}}}}{{{\rm{Total number of}}\,{\rm{N - O bond groups}}}}\\\\{\rm{Bond}}\,{\rm{order}}\,{\rm{ = }}\,\frac{4}{3}\\\\\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,{\rm{ = }}\,\,1.3.\\\end{array}

Therefore, nitrogen-to-oxygen bond order of the NO3{\rm{NO}}_3^ - molecule = 1.33.

The structure of NO3{\rm{NO}}_3^ - molecule is given below:

NO,

From the given structure of NO2{\rm{NO}}_2^ - molecule,

The total number of electron pairs in NO{\rm{N}} - {\rm{O}} bond = 3 electron pairs.

The total number of NO{\rm{N}} - {\rm{O}} bond groups = 2 bonds (1 double bond, 1 single bond).

The bond order is calculated below:

Bondorder=TotalnumberofelectronpairsinNObondTotalnumberofNObondgroupsBondorder=32=1.5\begin{array}{c}\\{\rm{Bond}}\,{\rm{order}}\,{\rm{ = }}\,\frac{{{\rm{Total number of electron pairs in N - O bond}}}}{{{\rm{Total number of}}\,{\rm{N - O bond groups}}}}\\\\{\rm{Bond}}\,{\rm{order}}\,{\rm{ = }}\,\frac{3}{2}\\\\\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,{\rm{ = }}\,\,1.5\\\end{array}

Therefore, nitrogen-to-oxygen bond order of the NO2{\rm{NO}}_2^ - molecule = 1.5

The structure of NO3{\rm{NO}}_3^ - molecule is given below:

NO, +
:0 :
x

From the given structure of NO2+{\rm{NO}}_2^ + molecule,

The total number of electron pairs in NO{\rm{N}} - {\rm{O}} bond = 4 electron pairs

The total number of NO{\rm{N}} - {\rm{O}} bond groups = 2 bonds(2 double bonds)

The bond order is calculated below:

Bondorder=TotalnumberofelectronpairsinNObondTotalnumberofNObondgroupsBondorder=42=2.\begin{array}{c}\\{\rm{Bond}}\,{\rm{order}}\,{\rm{ = }}\,\frac{{{\rm{Total number of electron pairs in N - O bond}}}}{{{\rm{Total number of}}\,{\rm{N - O bond groups}}}}\\\\{\rm{Bond}}\,{\rm{order}}\,{\rm{ = }}\,\frac{4}{2}\\\\\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,{\rm{ = }}\,\,2.\\\end{array}

Therefore, nitrogen-to-oxygen bond order in NO2{\rm{NO}}_2^ - molecule = 2

Ans:

The nitrogen-to-oxygen bond order is given below:

NO3
Toi
1.33
o 1.5
© 1.67
TO 2

The nitrogen-to-oxygen bond order is given below:

O
o
o
NO2
1
1.33
1.5
1.67

The nitrogen-to-oxygen bond order is given below:

*
0:
ta
NO2+
00000